OpenStudy (anonymous):

how many square meters of canvas is required for a conical tent whose height is 3.5 m and radius of whose base is 12m?(pie=22/7)

OpenStudy (anonymous):

in detail i want d answer

OpenStudy (anonymous):

?????????????????????????????

OpenStudy (anonymous):

area of canvas required = Curved Surface Area of the cone CSA of cone = π*r*l so we first need to find l l = sqroot(r^2 + h^2) = sqroot[(12m)^2 + (3.5m)^2] = sqroot[ 144 m^2 + 12.25 m^2]

OpenStudy (mihai):

471 square meters

OpenStudy (anonymous):

l = 12.5 m so CSA = 22/7 * 12m * 12.5 m = 471.43 sqmtr

OpenStudy (anonymous):

thxs to alllllllllllllllll

OpenStudy (anonymous):

answer is 471.4

OpenStudy (anonymous):

Hmm don't we consider the base?

OpenStudy (anonymous):

@vengeance921 we r not given that the base is also canvas. for the language above we just need to find the CSA

OpenStudy (anonymous):

True, thanks.

OpenStudy (anonymous):

Thanks, except for adding the base your calc was also right

OpenStudy (anonymous):

u hav done dat knw and u deleted it

OpenStudy (anonymous):

Yeah because I felt it's wrong so..